# Augmenting Smart Contract Decompiler Output through Fine-grained Dependency Analysis and LLM-facilitated Semantic Recovery

### v1

@@ 7. Discussion and Limitation
SmartHalo shares the following advantages in decompiler output optimization: (1) As evidenced by evaluation, SmartHalo is obviously effective in optimizing method boundaries, variable types and contract attributes, demonstrating superior performance compared to previous approaches; (2) SmartHalo establishes a novel combination of static analysis (SA) and large language models (LLMs), enabling the joint optimization on the decompiler output with the advantages of both SA and LLMs, thereby overcoming the limitation of prior works; (3) SmartHalo implements a rigorous correctness verification for the optimization output to eliminate the optimization errors. With the above-illustrated advantages, SmartHalo can optimize decompiler output precisely and comprehensively. All of the developers, participants, and third-party authorities can utilize SmartHalo to facilitate a number of downstream tasks in program comprehension and vulnerability analysis.
Threats to validity.  An external threat to validity is that the number of smart contract methods used in our evaluation here is relatively small (i.e., 456 smart contract methods). However, establishing the evaluation dataset and its ground-truth requires a large manual effort and incurs expensive cost of GPT fees. In addition, the effectiveness of SmartHalo remains unaffected by the limitation of dataset size, because our manual investigation confirms that the dataset involves diverse and main application scenarios of smart contract [^3] Note that the manual investigation results can be available via Section Data Availability, such as decentralized governance voting, token implementation, cross-chain protocols, and gambling.
An internal threat to validity is the potential bias in annotating the ground-truth in the 456 smart contract method of the evaluation dataset. We have mitigated this threat by inviting three domain researchers performing the annotation independently, and annotated the method boundaries, variable types and contract attributes via a manner of majority voting. Another internal threat to validity is the predefined static rules employed by SmartHalo for correctness verification (Section 4.3). However, SmartHalo remains a highly generalizable framework for correctness verification, because the predefined static rules are derived from the syntax of smart contracts, as outlined in Solidity documentation (Sol, [n. d.]). Hence, these static rules are essential and universally applicable across all Solidity contracts. Note that SmartHalo is incapable of supporting analyzing non-Solidity smart contract, because the decompiled code generated by its core decompiler Gigahorse (Grech et al., 2019) is specifically written in Solidity. However, SmartHalo offers the flexibility to integrate alternative decompilers tailored to other languages to extend its capabilities in supporting smart contracts written in other languages. Importantly, the effectiveness of SmartHalo is still guaranteed by the fact that most smart contracts are written in Solidity language.
8. Related Work
@@
Smart Contract Decompilation.  The decompilation of smart contracts has garnered significant attention due to the prevalence of bytecode-only smart contracts (Suiche, [n. d.]; Brent et al., 2018; Liu et al., 2023). Gigahorse (Grech et al., 2019) translates smart contracts from low-level EVM bytecode into a 3-address code representation. Elipmoc (Grech et al., 2022) employs a novel context sensitivity called transactional sensitivity to achieve a more effective static abstraction. Erays (Zhou et al., 2018) and EtherSolve (Contro et al., 2021) aim to recover the CFG from the EVM bytecode. Recently, SigRec (Chen et al., 2022b) has been proposed to recover the signatures of public functions. DeepInfer (Zhao et al., 2023) leverages deep learning techniques to infer function signatures.
@@
Decompilation Optimization.  The existing decompilation optimization research mainly focus on other languages such as C++ and Java (Pei et al., 2020), which can be divided into the following two categories (Pang et al., 2021; Haq and Caballero, 2021). The first category aims at recovering the semantic information from the assembly code or intermediate language. Among them, previous works, including DEBIN (He et al., 2018), TIE (Lee et al., 2011), OSPREY (Zhang et al., 2021), and BDA (Zhang et al., 2019), analyze the program dependency via the static analysis. However, they are with low coverage due to their reliance on heuristic rules. Another group of works, including Nero (David et al., 2020), NFRE (Gao et al., 2021), SYMLM (Jin et al., 2022), focus on designing encoder-decoder architecture models to predict the method name for the binary. The other category of work focuses on optimizing variable name, variable type, and structure of code generated by decompiler. For example, Jaffe et al. utilize statistic machine translation to recover variable names (Jaffe et al., 2018). DIRE (Lacomis et al., 2019) leverages neural machine translation to predict variable names. DIRTY (Chen et al., 2022a) with the transformer-based model to predict the variable types and names. DeGPT (Hu et al., 2024) applies the LLM to augment the decompiler output. However, these works ignore the static domain knowledge, and are prone to making prediction errors.
9. Conclusion
@@
In this paper, we introduced SmartHalo, a novel framework designed to address critical deficiencies in existing Solidity smart contract decompilers. By leveraging the strengths of static analysis (SA) and large language models (LLM), SmartHalo significantly enhances the accuracy and reliability of decompiled code. Our framework systematically recovers method boundaries, variable types, and contract attributes, facilitating the program comprehension and vulnerability detection. Extensive evaluation demonstrates that SmartHalo outperforms state-of-the-art decompilers, showing substantial improvements in method boundary identification, variable type inference, and contract attribute recovery. We also showed that the output of SmartHalo can significantly enhance the effectiveness of downstream tasks for vulnerability detection.
Data Availability
Currently, the artifact of SmartHalo, corresponding evaluation dataset, and other online materials are available in this anonymous repository (https://figshare.com/s/5d4fc1ad2312a4be9370).

### v2

@@ VII Discussion and Limitation
SmartHalo shares the following advantages in decompiler output optimization: (1) As evidenced by evaluation, SmartHalo is obviously effective in optimizing function boundaries, variable types and contract attributes, demonstrating superior performance compared to previous approaches; (2) SmartHalo establishes a novel combination of static analysis (SA) and large language models (LLMs), enabling the joint optimization on the decompiler output with the advantages of both SA and LLMs, thereby overcoming the limitation of prior works. (3) SmartHalo implements a rigorous correctness verification for the optimization output to eliminate the optimization errors.
SmartHalo has good adaptability to newly-deployed contracts due to its flexible and generic core components: (1) Static analysis employs typical control‑flow analysis to extract the dependencies that are generic to all contracts. These dependency‑extraction methods are equally applicable to newly deployed contracts; (2) The LLMs is adept at few-shot in-context learning, which refers to the capability of LLMs to instantaneously comprehend and address new and unseen cases/tasks during the inference stage, based solely on a small number of known examples (shots). This helps SmartHalo to be feasible to the newly-deployed contracts.
Although our evaluation of SmartHalo focuses on the function level, SmartHalo is actually capable of supporting contract‑level optimizations. Specifically, the prompts generated by static analysis embed complete code-context information which captures all the contents relevant to the optimization targets across the whole contract, which enables SmartHalo to perform contract-level optimization. For example, the prompt templates of Table I introduce the contextual information of modifiers (i.e., row 10), which guides the LLM to reconstruct the corresponding modifier functions; and as highlighted by the yellow part of Fig. 6(a), the COT prompts produced by static analysis effectively capture the dependencies between key–value pairs within complex storage patterns (e.g., mappings and structs), this capability facilitates the recovery of the complex storage variable types ( i.e., mapping (bytes 32 = uint256) data). Further, as evidenced by Section V-E, we have evaluated SmartHalo’s feasibility to complex and diverse contracts. And the results confirm that SmartHalo exhibits good scalability for contract-level decompilation optimization.
The evolution of LLM APIs presents potential challenges to the reproducibility of evaluation results. To address this concern, we have implemented the following strategies. Firstly, we explicitly list the exact API configurations for each model in the SmartHalo’s repository. These details encompass the server provider, model identifier, version/release date, invocation endpoints, context length limitations, and relevant environment variables. Further, for open-source models, except for API configurations, we provide an access link of the repository corresponding to each model version used in this paper. Finally, as evidenced by Section V-E, SmartHalo maintains a good performance on the evaluated smaller language model. In future work, we can explore integrating more alternative fine-tuned smaller language models into SmartHalo to further reduce the effects of evolving LLMs on reproducibility.
A current limitation is that SmartHalo cannot recover inheritance structures within contracts. The reason for this limitation is that, inheritance relationship and class information are missed at the bytecode level in smart contracts. Particularly, state-of-the-art efforts [9], [19] in other domains (e.g., C++, Java) also cannot recover the inheritance structure for smart contract, because they mainly rely on class information. Given the difficulty of this task, we leave the exploration of inheritance structure recovery to future work.
Threats to validity.  An external threat to validity is that the number of smart contract functions used in our evaluation here is relatively small (i.e., 456 smart contract functions). However, we consider the dataset size sufficient for the following reasons. Firstly, the dataset was randomly selected from the largest real-world contract repository. Secondly, the dataset size is similar to those used in SOTA studies [4, 18]. Thirdly, our manual investigation confirms that the dataset involves diverse and mainstream application scenarios of smart contract [^2] Note that the manual investigation results are available via https://figshare.com/s/5d4fc1ad2312a4be9370?file=49346911.
An internal threat to validity is the predefined type rules employed by SmartHalo for correctness verification (Section IV-C). SmartHalo remains a highly generalizable framework for the following reasons: First, SmartHalo focuses on optimizing Solidity decompiler output. This setup is aligned with recent contract analysis research [1, 44]. Second, most contracts are written in Solidity [1] and are widely deployed on major blockchains (e.g., Ethereum, TRON, and BNB). Third, the static rules(see Figure 7) are generic across different Solidity versions, as they reflect knowledge of fundamental types (e.g.,array). A review of Solidity’s release-changelogs indicates that these rules remained unchanged across all Solidity’s major updates [45]. Fourth, SmartHalo is backward-compatible, can easily adapt to contract evolution with corresponding knowledge.
VIII Related Work
@@
Smart Contract Decompilation.  The decompilation of smart contracts has garnered significant attention due to the prevalence of bytecode-only smart contracts [46] Gigahorse [9] translates smart contracts from low-level EVM bytecode into a 3-address code representation. Elipmoc [38] is the 2.0 verson of Gigahorse, which employs a novel context sensitivity called transactional sensitivity to achieve a more effective static abstraction. Erays [47] and EtherSolve [48] aim to recover the CFG from the EVM bytecode. Recently, SigRec [12] has been proposed to recover the signatures of public functions. DeepInfer [11] leverages deep learning techniques to infer function signatures.
@@
Decompilation Optimization.  The existing decompilation optimization research mainly focus on other languages such as C++ and Java [49, 50], which can be divided into the following two categories [51]. The first category aims at recovering the semantic information from the assembly code or intermediate language. Among them, previous works, including DEBIN [15], OSPREY [52], and BDA [53], analyze the program dependency via the static analysis. However, they are with low coverage due to their reliance on heuristic rules. Another group of works, including Nero [54], NFRE [55], SYMLM [56], focus on designing encoder-decoder architecture models to predict the function name for the binary. The other category of work focuses on optimizing variable name, variable type, and structure of code generated by decompiler, including DIRE [57], DIRTY [16] and DeGPT [4].
IX Conclusion
@@
In this paper, we introduced SmartHalo, a novel framework designed to address critical deficiencies in existing Solidity smart contract decompilers. By leveraging the strengths of static analysis (SA) and large language models (LLMs), SmartHalo significantly enhances the accuracy and readability of decompiled code. Extensive evaluation demonstrates that SmartHalo outperforms state-of-the-art decompilers, showing substantial improvements in function boundary identification, variable type inference, and contract attribute recovery. We also showed that the output of SmartHalo can significantly enhance the effectiveness of downstream tasks.